calculate the new freezing point of water when 72.0 g of BaCl2 are added to 0.175 kg of water


1
Expert's answer
2022-05-26T13:03:03-0400

1. Find the molality of the solution.

molality is the ratio of the amount of solute (mol) to the mass of solvent (kg):

m= n(BaCl2) / 0,175kg = 0,346/0,175 =1,977


[ n(BaCl2) = 72 / 208 = 0,346 mol ]


2. Now substitute this value into the freezing point expression and solve for ∆T:

∆T = i * m * Kf

i – van't Hoff Factor= 3 (BaCl2 =>Ba2+ + 2Cl-

dissociation produces 3 ions);

Kf — freezing point constant = 1.86

∆T =imKf = 3 * 1,977 * 1.86 = 11,03° C


New freezing point = 0 - 11,03 = –11,03° C

ANSWER: –11,03° C